Can you help? This take home exercise has me stumped:
Write the quadratic equation, given the solution: (3/5, 2/7)
Thank you.

Answer by edjones(8007)

Do things backwards.
x=3/5, x=2/7
(x-(3/5))(x-(2/7))=0
x^2-(3x/5)-(2x/7)+(6/35)=0
-(3/5)-(2/7)=-(21/35)-(10/35)=-31/35
x^2-(31x/35)+(6/35)=0
Ed